OpenStudy (anonymous):

Find g'(x) when g(x)= ln{(x-1)/(x+1)}

OpenStudy (anonymous):

\[\ln \frac{x-1}{x+1}=\ln (x-1)-\ln(x+1) \ \ \]

OpenStudy (anonymous):

\[\frac{d}{dx}\ln f(x)=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

Could you use the hints to find your answer?

OpenStudy (anonymous):

Not yet still tryin

OpenStudy (anonymous):

Oh damn I forgot square root everything inside brackets